# Fully Dynamic Euclidean $k$-Means

### v2

@@ 10.3 Application: Maintaining Cluster Weight
Before presenting our construction for Lemma 10.14, we state the following corollary, which follows as a consequence of Lemmas 10.14 and 9.4, and will be useful in the analysis of restricted $k$-means (Section 11).
@@
There exists a data structure that maintains an ordering $c_{1},c_{2},\ldots,c_{|S|}$ of centers $S$ such that for each $11\leq i\leq|S|-$ 1, the following holds,
@@
Since both the values $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$ and the weights $\{w_{S}(s)\}_{s\in S}$ are explicitly maintained by Lemmas 9.4 and 10.14, respectively, the desirable ordering $c_{1},c_{2},\ldots,c_{|S|}$ of the centers $S$ of centers $S$ can be maintained synchronously, with only a $\tilde{O}(1)$-factor increase in update time. ∎
@@
  • $w_{S}(s):=\sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(i,z)+w(s)$,141414 If $s\notin X$, we define $w(s)=0$. for every center $s\in S$.
@@
$w_{S}(s)$ $\leavevmode\nobreak\ \leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(s)+w(s)$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w(\operatorname{bk}(i,z))+w(s)$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}\sum_{x\in\operatorname{bk}(i,z)}w(s)+w(s).$
Since $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of $X-S$ (Claim 10.5) and $\sigma(x)=\sigma^{\prime}(i,z)$ for every $(i,z)\in H^{\prime}$ and every $x\in\operatorname{bk}(i,z)$ (Claim 10.7). It then follows that
$w_{S}(s)\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{x\in X-S:\sigma(x)=s}w(x)+w(s)=\sum_{x\in X:\sigma(x)=s}w(x).$
@@ 10.4 Application: $D^{2}$-Sampling
$\forall x\in X:$ $\Pr[y=x]\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\epsilon^{-6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$ (24)
@@
$y\leavevmode\nobreak\ \in\leavevmode\nobreak\ \bigcup_{(i,z)\in H^{\prime}}\bigcup_{z^{\prime}\in f(i,z)}\operatorname{bk}(i,z,z^{\prime})\leavevmode\nobreak\ =\leavevmode\nobreak\ \bigcup_{(i,z)\in H^{\prime}}\operatorname{bk}(i,z)\leavevmode\nobreak\ =\leavevmode\nobreak\ X-S.$ (25)
Consider a specific point $x\in X$. If $x\in S$ is a center, then by Equation 25, we have $\Pr[y=x]=0$. Since $\mathrm{dist}(x,S)=0$, Equation 24 trivially holds in this case. If $x\notin S$ is not a center, then there exists a unique pair $(i_{x},z_{x})\in H^{\prime}$ such that $x\in\operatorname{bk}(i_{x},z_{x})=\bigcup_{z^{\prime}\in f(i_{x},z_{x})}\operatorname{bk}(i,z,z^{\prime})$, since the collection $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ forms a partition of $X-S$ (Claim 10.5). Moreover, since the low-level buckets $\{\operatorname{bk}(i_{x},z_{x},z^{\prime})\}_{z^{\prime}\in f(i_{x},z_{x})}$ are disjoint (Item 4 of Definition 10.1), there exists a unique value $z_{x}^{\prime}\in f(i_{x},z_{x})$ such that $x\in\operatorname{bk}(i_{x},z_{x},z_{x}^{\prime})$. Thus, the point $x$ is sampled with probability
$\Pr[y=x]$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{w_{H}(i_{x},z_{x})}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}\cdot\frac{w_{L}(i_{x},z_{x},z^{\prime}_{x})}{\sum_{z^{\prime}\in f(i_{x},z_{x})}w_{L}(i_{x},z_{x},z^{\prime})}\cdot\frac{w(x)}{w(\operatorname{bk}(i_{x},z_{x},z^{\prime}_{x}))}$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \frac{(3\Gamma)^{2i_{x}}\cdot w(x)}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}.$
@@
$\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}(3\Gamma)^{2i}\cdot\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}w(\operatorname{bk}(i,z))\cdot(3\Gamma)^{2i}$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}w(x)\cdot(3\Gamma)^{2i}$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}O(\Gamma^{4})\cdot w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 10.6)
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\Gamma^{4})\cdot\sum_{x\in X-S}w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 10.5)
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\Gamma^{4})\cdot\textsc{cost}(X,S).$
@@
$\Pr[y=x]\leavevmode\nobreak\ \geq\leavevmode\nobreak\ \Omega(\Gamma^{4})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}\leavevmode\nobreak\ =\leavevmode\nobreak\ \Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$
@@ 11 Restricted $k$-Means
Assuming that the weighted point set $X$ and the center set $S$ are maintained by the data structures previously discussed (specifically Corollaries 10.15, 9.2 and 10.14), we design a subroutine that computes an approximate restricted $k$-means solution in sublinear time.
@@
Given access to the data structures from Corollaries 10.15, 9.2 and 10.14, there exists a subroutine that, on input an integer $1\leq r\leq|S|$, computes a size-$r$ subset $R\subseteq S$ in time $\tilde{O}(2^{\epsilon d}r+r^{1+\epsilon})$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
$\textsc{cost}(X,S-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
@@
Recall that all data structures from Corollaries 10.15, 9.2 and 10.14 support dynamic updates to both $X$ and $S$ against an adaptive adversary, each with update time $\tilde{O}(2^{\epsilon d})$. By packing them into a single data structure together with the subroutine in Lemma 11.1, we obtain Lemma 7.5 stated in Section 7.
@@
3:Compute a size-$r$ subset $\widehat{R}\subseteq T:=T_{1}+T_{2}$ such that $\triangleright$ $T\subseteq S$.
$\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-\widehat{R})\leq O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-R),$
where $w_{S}:S\to\mathbb{R}_{\geq 0}$ is the weight function maintained by Lemma 10.14.
4:\$\widehat{R}$.
@@
3 can be implemented in time $\tilde{O}(r^{1+\epsilon})$: Since the weight function $w_{S}:S\to\mathbb{R}_{\geq 0}$ is explicitly maintained by Lemma 10.14, we can construct the weighted set $T=T_{1}\cup T_{2}$, where every point $s\in T$ has weight $w_{S}(s)$, in time $\tilde{O}(r)$. Then, to compute $\widehat{R}$, it suffices to run an $(|T|-r)$-means algorithm on the weighted set $T$, with centers selected from $T$, and obtain a solution $\widehat{S}\subseteq T$ with $|\widehat{S}|=|T|-r$. Then, we return $\widehat{R}\leftarrow T-\widehat{S}$ as the desired output; for this, we can apply the static algorithm from [JJLL25], which takes $\tilde{O}(r^{1+\epsilon})$ time.
@@
$\textsc{cost}(X,S-\widehat{R})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$ (26)
@@
$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$
@@
$\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R).$ (27)
@@
$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
@@
$\textsc{cost}(T,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
We defer the proof of Claims 11.3 and 11.4 to Sections 11.2 and 11.4, respectively.
@@
LHS of Equation 27 $=\leavevmode\nobreak\ \min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$
$\leq\leavevmode\nobreak\ \textsc{cost}(T,T-\widehat{R}^{*})$ ($\widehat{R}^{*}\subseteq T_{1}\subseteq T$, $|\widehat{R}^{*}|=r$)
$\leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-\widehat{R}^{*})$ (Claim 11.4)
$\leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\textsc{cost}(S,S-R^{*})$ (Claim 11.3)
$=\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Definition of $R^{*}$)
@@
$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}(s,T-R)\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$ (28)
@@
$\textsc{cost}(X,S-\widehat{R})$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot\mathrm{dist}^{2}(x,S-\widehat{R})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot(\mathrm{dist}(x,\sigma(x))+\mathrm{dist}(\sigma(x),S-\widehat{R}))^{2}$ (triangle inequality)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(\sigma(x),S-\widehat{R}))$ (AM-GM inequality)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\textsc{cost}(X,S)+2\cdot\sum_{s\in S}w_{S}(x)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$ (Properties (i) and (ii))
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)+2\cdot\textsc{cost}(S,S-\widehat{R}).$
The last step applies $\textsc{cost}(X,S)\leq\mathrm{OPT}_{|S|-r}^{S}(X)$ and $\textsc{cost}(S,S-\widehat{R})=\sum_{s\in S}w_{S}(x)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$. Since every center $s\in S-\widehat{R}$ has a zero distance to $S-\widehat{R}$, we have
$\textsc{cost}(S,S-\widehat{R})$ $\leavevmode\nobreak\ =\leavevmode\nobreak\ \sum_{s\in\widehat{R}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in\widehat{R}}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-\widehat{R})$ ($T\subseteq S$)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ \sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-\widehat{R})$ ($\widehat{R}\subseteq T$)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$ (3 of Algorithm 4)
@@
$\textsc{cost}(S,S-\widehat{R})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Equation 28)
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\textsc{cost}(S,S-R^{*}_{X})$ ($R^{*}_{X}\subseteq S$, $|R^{*}_{X}|=r$)
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,s)+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
$\leavevmode\nobreak\ =\leavevmode\nobreak\ O(\epsilon^{-7})\cdot\sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\textsc{cost}(X,S)+O(\epsilon^{-7})\cdot\textsc{cost}(X,S-R^{*}_{X})$ (Property (i))
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
@@
$\textsc{cost}(X,S-\widehat{R})$ $\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-6})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)+O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X)$
$\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
@@
$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
@@
$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak\ \leq\leavevmode\nobreak\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$ (29)
@@
$\textsc{cost}(S,S-\widehat{R}^{*})\leavevmode\nobreak$ $=\leavevmode\nobreak\ \sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})$

### v3

@@ 11.3 Application: Maintaining Cluster Weight
Before presenting our construction for Lemma 11.14, we state the following corollary, which follows as a consequence of Lemmas 11.14 and 10.4, and will be useful in the analysis of restricted $k$-means (Section 12).
@@
There exists a data structure that maintains an ordering $c_{1},c_{2},\ldots,c_{|S|}$ of centers $S$ such that for each $1\leq i\leq|S|-1$, the following holds,
@@
Since both the values $\{\widehat{\mathrm{dist}}(s,S-s)\}_{s\in S}$ and the weights $\{w_{S}(s)\}_{s\in S}$ are explicitly maintained by Lemmas 10.4 and 11.14, respectively, the desirable ordering $c_{1},c_{2},\ldots,c_{|S|}$ of the centers $S$ of centers $S$ can be maintained synchronously, with only a $\tilde{O}(1)$-factor increase in update time. ∎
@@
  • $w_{S}(s):=\sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(i,z)+w(s)$, [^14] If $s\notin X$, we define $w(s)=0$. for every center $s\in S$.
@@
$w_{S}(s)$ $\penalty 10000\ \penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w_{H}(s)+w(s)$
$\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}w(\operatorname{bk}(i,z))+w(s)$
$\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}:\sigma^{\prime}(i,z)=s}\sum_{x\in\operatorname{bk}(i,z)}w(s)+w(s).$
Since $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ is a partition of XSX-S (Claim 11.5) and $\sigma(x)=\sigma^{\prime}(i,z)$ for every $(i,z)\in H^{\prime}$ and every $x\in\operatorname{bk}(i,z)$ (Claim 11.7). It then follows that
$w_{S}(s)\penalty 10000\ =\penalty 10000\ \sum_{x\in X-S:\sigma(x)=s}w(x)+w(s)=\sum_{x\in X:\sigma(x)=s}w(x).$
@@ 11.4 Application: $D^{2}$-Sampling
$\forall x\in X:$ $\Pr[y=x]\penalty 10000\ \geq\penalty 10000\ \Omega(\epsilon^{-6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$ (25)
@@
$y\penalty 10000\ \in\penalty 10000\ \bigcup_{(i,z)\in H^{\prime}}\bigcup_{z^{\prime}\in f(i,z)}\operatorname{bk}(i,z,z^{\prime})\penalty 10000\ =\penalty 10000\ \bigcup_{(i,z)\in H^{\prime}}\operatorname{bk}(i,z)\penalty 10000\ =\penalty 10000\ X-S.$ (26)
Consider a specific point $x\in X$. If $x\in S$ is a center, then by Equation 26, we have $\Pr[y=x]=0$. Since $\mathrm{dist}(x,S)=0$, Equation 25 trivially holds in this case. If $x\notin S$ is not a center, then there exists a unique pair $(i_{x},z_{x})\in H^{\prime}$ such that $x\in\operatorname{bk}(i_{x},z_{x})=\bigcup_{z^{\prime}\in f(i_{x},z_{x})}\operatorname{bk}(i,z,z^{\prime})$, since the collection $\{\operatorname{bk}(i,z)\}_{(i,z)\in H^{\prime}}$ forms a partition of XSX-S (Claim 11.5). Moreover, since the low-level buckets $\{\operatorname{bk}(i_{x},z_{x},z^{\prime})\}_{z^{\prime}\in f(i_{x},z_{x})}$ are disjoint (Item 4 of Definition 11.1), there exists a unique value $z_{x}^{\prime}\in f(i_{x},z_{x})$ such that $x\in\operatorname{bk}(i_{x},z_{x},z_{x}^{\prime})$. Thus, the point $x$ is sampled with probability
$\Pr[y=x]$ $\penalty 10000\ =\penalty 10000\ \frac{w_{H}(i_{x},z_{x})}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}\cdot\frac{w_{L}(i_{x},z_{x},z^{\prime}_{x})}{\sum_{z^{\prime}\in f(i_{x},z_{x})}w_{L}(i_{x},z_{x},z^{\prime})}\cdot\frac{w(x)}{w(\operatorname{bk}(i_{x},z_{x},z^{\prime}_{x}))}$
$\penalty 10000\ =\penalty 10000\ \frac{(3\Gamma)^{2i_{x}}\cdot w(x)}{\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)}.$
@@
$\sum_{(i,z)\in H^{\prime}}w_{H}(i,z)$ $\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}}(3\Gamma)^{2i}\cdot\sum_{z^{\prime}\in f(i,z)}w(\operatorname{bk}(i,z,z^{\prime}))$
$\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}}w(\operatorname{bk}(i,z))\cdot(3\Gamma)^{2i}$
$\penalty 10000\ =\penalty 10000\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}w(x)\cdot(3\Gamma)^{2i}$
$\penalty 10000\ \leq\penalty 10000\ \sum_{(i,z)\in H^{\prime}}\sum_{x\in\operatorname{bk}(i,z)}O(\Gamma^{4})\cdot w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 11.6)
$\penalty 10000\ =\penalty 10000\ O(\Gamma^{4})\cdot\sum_{x\in X-S}w(x)\cdot\mathrm{dist}^{2}(x,S)$ (Claim 11.5)
$\penalty 10000\ =\penalty 10000\ O(\Gamma^{4})\cdot\textsc{cost}(X,S).$
@@
$\Pr[y=x]\penalty 10000\ \geq\penalty 10000\ \Omega(\Gamma^{4})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}\penalty 10000\ =\penalty 10000\ \Omega(\epsilon^{6})\cdot\frac{w(x)\cdot\mathrm{dist}^{2}(x,S)}{\textsc{cost}(X,S)}.$
@@ 12 Restricted $k$-Means
Assuming that the weighted point set $X$ and the center set $S$ are maintained by the data structures previously discussed (specifically Corollaries 11.15, 10.2 and 11.14), we design a subroutine that computes an approximate restricted $k$-means solution in sublinear time.
@@
Given access to the data structures from Corollaries 11.15, 10.2 and 11.14, there exists a subroutine that, on input an integer $1\leq r\leq|S|$, computes a size-$r$ subset $R\subseteq S$ in time $\tilde{O}(2^{\epsilon d}r+r^{1+\epsilon})$, such that the following holds with probability $1-\frac{1}{\mathrm{poly}(\Delta^{d})}$:
$\textsc{cost}(X,S-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X),$
@@
Recall that all data structures from Corollaries 11.15, 10.2 and 11.14 support dynamic updates to both $X$ and $S$ against an adaptive adversary, each with update time $\tilde{O}(2^{\epsilon d})$. By packing them into a single data structure together with the subroutine in Lemma 12.1, we obtain Lemma 8.5 stated in Section 8.
@@
3: Compute a size-$r$ subset $\widehat{R}\subseteq T:=T_{1}+T_{2}$ such that $\triangleright$ $T\subseteq S$. $\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-\widehat{R})\leq O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\sum_{x\in T}w_{S}(x)\cdot\mathrm{dist}^{2}(x,T-R),$ where $w_{S}:S\to\mathbb{R}_{\geq 0}$ is the weight function maintained by Lemma 11.14.
4: return $\widehat{R}$.
@@
3 can be implemented in time $\tilde{O}(r^{1+\epsilon})$: Since the weight function $w_{S}:S\to\mathbb{R}_{\geq 0}$ is explicitly maintained by Lemma 11.14, we can construct the weighted set $T=T_{1}\cup T_{2}$, where every point $s\in T$ has weight $w_{S}(s)$, in time $\tilde{O}(r)$. Then, to compute $\widehat{R}$, it suffices to run an (|T|r)(|T|-r)-means algorithm on the weighted set $T$, with centers selected from $T$, and obtain a solution $\widehat{S}\subseteq T$ with $|\widehat{S}|=|T|-r$. Then, we return $\widehat{R}\leftarrow T-\widehat{S}$ as the desired output; for this, we can apply the static algorithm from [abs-2504-03513], which takes $\tilde{O}(r^{1+\epsilon})$ time.
@@
$\textsc{cost}(X,S-\widehat{R})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$ (27)
@@
$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}^{2}(s,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$
@@
$\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R).$ (28)
@@
$\textsc{cost}(S,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
@@
$\textsc{cost}(T,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R).$
We defer the proof of Claims 12.3 and 12.4 to Sections 12.2 and 12.4, respectively.
@@
LHS of Equation 28 $=\penalty 10000\ \min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$
$\leq\penalty 10000\ \textsc{cost}(T,T-\widehat{R}^{*})$ ($\widehat{R}^{*}\subseteq T_{1}\subseteq T$, $|\widehat{R}^{*}|=r$)
$\leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-\widehat{R}^{*})$ (Claim 12.4)
$\leq\penalty 10000\ O(\epsilon^{-6})\cdot\textsc{cost}(S,S-R^{*})$ (Claim 12.3)
$=\penalty 10000\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Definition of $R^{*}$)
@@
$\min_{R\subseteq T:|R|=r}\sum_{s\in T}w_{S}(s)\cdot\mathrm{dist}(s,T-R)\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-6})\cdot\min_{R\subseteq S:|R|=r}\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R).$ (29)
@@
[…]
$\penalty 10000\ \leq\penalty 10000\ \sum_{x\in X}w(x)\cdot(\mathrm{dist}(x,\sigma(x))+\mathrm{dist}(\sigma(x),S-\widehat{R}))^{2}$ (triangle inequality)
$)\displaystyle\penalty 10000\ \leq\penalty 10000\ \sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(\sigma(x),S-\widehat{R}$)) (AM-GM inequality)
[…]
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-1})\cdot\min_{R\subseteq T:|R|=r}\textsc{cost}(T,T-R)$ (3 of Algorithm 4)
@@
[…]
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-7})\cdot\min_{R\subseteq S:|R|=r}\textsc{cost}(S,S-R)$ (Equation 29)
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-7})\cdot\textsc{cost}(S,S-R^{*}_{X})$ ($R^{*}_{X}\subseteq S$, $|R^{*}_{X}|=r$)
$\penalty 10000\ =\penalty 10000\ O(\epsilon^{-7})\cdot\sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
$\penalty 10000\ =\penalty 10000\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot\mathrm{dist}^{2}(s,S-R^{*}_{X})$
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-7})\cdot\sum_{s\in S}\sum_{x\in\sigma^{-1}(s)}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,s)+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
$\penalty 10000\ =\penalty 10000\ O(\epsilon^{-7})\cdot\sum_{x\in X}w(x)\cdot 2\cdot(\mathrm{dist}^{2}(x,\sigma(x))+\mathrm{dist}^{2}(x,S-R^{*}_{X}))$
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\textsc{cost}(X,S)+O(\epsilon^{-7})\cdot\textsc{cost}(X,S-R^{*}_{X})$ (Property (i))
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
@@
[…]
$\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-13})\cdot\mathrm{OPT}_{|S|-r}^{S}(X).$
@@
$\textsc{cost}(S,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$
@@
$\textsc{cost}(S,S-\widehat{R}^{*})\penalty 10000\ \leq\penalty 10000\ O(\epsilon^{-3})\cdot\textsc{cost}(S,S-R^{*}).$ (30)
@@
$\textsc{cost}(S,S-\widehat{R}^{*})\penalty 10000\$$=\penalty 10000\ \sum_{s\in S}w_{S}(s)\cdot\mathrm{dist}^{2}(s,S-\widehat{R}^{*})$